OpenStudy (anonymous):

I need to solve the equation. 9|9 – 8x| = 2x + 3

OpenStudy (anonymous):

x=6/5

OpenStudy (anonymous):

check for yourself its right hun !

OpenStudy (anonymous):

I need the work for the problem too

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

x=39/37 = x=6/5

OpenStudy (anonymous):

How do you get that answer?

OpenStudy (anonymous):

My calculator :]

OpenStudy (anonymous):

But, what are the steps? Haha

OpenStudy (anonymous):

Oh lol idk my calculator just solved it . Sorry :/

OpenStudy (anonymous):

Well, i have no clue how to do it.

OpenStudy (anonymous):

two solutions 6/5 and 39/37 1 sec writting how to solve :x

OpenStudy (dumbcow):

always get the absolute value part alone on 1 side first |9-8x| = (2x+3)/9 then split it into 2 cases, pos and neg --> 9-8x = (2x+3)/9 and --> 9-8x = -(2x+3)/9 solve each equation for x

OpenStudy (anonymous):

|9-8x| -> (9-8x) and (8x-9) 9 (9 - 8x) = 2x + 3 81 - 72x = 2x + 3 78 = 74x x = 78/74 = 39/37 9 (8x - 9) = 2x + 3 72x - 81 = 2x + 3 70x = 84 x = 84/70 = 6/5

OpenStudy (anonymous):

THANK YOU GUYS FOR EXPLAINING IT =]

OpenStudy (anonymous):

Thanks so muchh!

OpenStudy (lgbasallote):

what bugs me though is...why write the answer *before* a solution? what's the point in that? o.O

OpenStudy (anonymous):

:c idk lol just wanted to correct babyboy's answer :s then proceeded to solvee